ACCOUNT Work in Process—Baking Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Dec. 1 Bal., 24,000 units, 3⁄4 completed 116,700
31 Direct materials, 134,000 units 234,500 351,200
31 Direct labor 150,000 501,200
31 Factory overhead 375,000 876,200
31 Goods finished, 128,000 units 760,700 115,500
31 Bal., ? units, 1⁄2 completed 115,500

a. Based on the above data, determine each cost listed below. Round "cost per equivalent unit" answers to the nearest cent.

1. Direct materials cost per equivalent unit $
2. Conversion cost per equivalent unit $
3. Cost of the beginning work in process completed during December $
4. Cost of units started and completed during December $
5. Cost of the ending work in process $

b. Assuming that the direct materials cost is the same for November and December, did the conversion cost per equivalent unit increase, decrease, or remain the same in December?
Answer #1

Ending work in process = 24000+134000-128000 = 30000

Equivalent unit of material = (24000*0%+104000*100%+30000*100%) = 134000

Equivalent unit of conversion = (24000*1/4+104000*100%+30000*1/2) = 125000

1) Direct material Cost per equivalent unit = 234500/134000 = 1.75

2) Conversion Cost per equivalent unit = (150000+375000)/125000 = 4.2

3) Cost of beginning work in process completed = (6000*4.2+116700) = 141900

4) Cost of units started and completed = 104000*5.95 = 618800

5) Ending work in process = (30000*1.75+15000*4.2) = 115500

b) Direct material cost in beginning work in process = 24000*1.75 = 42000

Conversion Cost in beginning work in process = 116700-42000 = 74700

Conversion Cost per equivalent unit in beginning WIP = 74700/18000 = 4.15

Conversion Cost per equivalent unit increase by (4.20-4.15) = 0.05 per unit